# Jacquet modules of Tate cohomology and base change lifting

### v1

@@ 3.5. Genericity of Tate cohomology
An equivalent condition for $(\pi,V)$ to be $\mathfrak{m}_{F}$-generic–which we will use frequently–is that the twisted Jacquet module $V_{U(F),\Theta_{F}}\simeq R$. We also have the obvious notions of $\mathfrak{m}_{E}$ generic representations of $G(E)$. Let $(\pi,V)$ be a smooth generic $R$-representation of $G(F)$. Let $\mathcal{J}_{\pi}$ be a non-zero linear functional in ${\rm Hom}_{U(F)}(V,\Theta_{F})$. Let $\mathbb{W}(\pi,\Theta_{F})\subset{\rm Ind}_{U(F)}^{G(F)}(\Theta_{F})$ be the space of functions $W_{v}:G(F)\rightarrow R$, $v\in V$, where
@@
3.5.2.
We now prove that the ${\rm Gal}(E/F)$ Tate cohomology of a generic $\ell$-modular ${\rm Gal}(E/F)$-stable representation of $G(E)$ has a unique generic sub-quotient. We proved this result for the case where $G={\rm GL}_{n}/F$ using Kirillov model. However, we give a simpler proof here. Which will be used in the later sections to prove irreducibility of Tate cohomology of certain representations of ${\rm GL}_{n}$ obtained as Base change lifts of cuspidal representations.
Theorem 3.5.
Let $E/F$ be a finite Galois extension with $[E:F]=\ell$. Let $G$ be a quasi-split reductive algebraic group defined over $F$. Let $(\Pi,V)$ be a ${\rm Gal}(E/F)$ stable $\mathfrak{m}_{E}$-generic $\ell$-modular representation of $G(E)$. Then the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\Pi)$ admits a unique $\mathfrak{m}_{F}$-generic sub-quotient.
@@
Let $\Gamma$ be the Galois group ${\rm Gal}(E/F)$. Let $\mathcal{J}$ be a non-zero element in ${\rm Hom}_{U(E)}(\Pi,\Theta_{E})$. Let $\mathcal{W}$ be the Whittaker model of $\Pi$. Using Proposition 3.2, the Tate cohomology applied to $\mathcal{W}\rightarrow\mathcal{W}_{U(E),\Theta_{E}}$ gives the following injection:
@@
Let $\sigma$ be a generator of ${\rm Gal}(E/F)$. Since ${\rm Gal}(E/F)$ acts trivially on the twisted Jacquet module $\mathcal{W}_{U(E),\Theta_{E}}\simeq\overline{\mathbb{F}}_{\ell}$, the Tate cohomology group $\widehat{H}^{0}(\mathcal{W}_{U(E),\Theta_{E}})$ is a one dimensional vector space over $\overline{\mathbb{F}}_{\ell}$. Now, let us consider the restriction to $G(F)$ map
@@
It is enough to show that the map $\Phi$ is non-zero. Suppose $\Phi(W_{v})=0$ for all $W_{v}\in\mathcal{W}^{\Gamma}$. Let $v\in{\rm Ker}(\mathcal{J})\cap\mathcal{W}^{\Gamma}$. Note that $\Pi(g)v\in\operatorname{Ker}(\mathcal{J})$, for $g\in G(F)$ and $g\in U(E)$. This shows that $\Pi(g)v\in{\rm Ker}(\mathcal{J})$, for all $g\in G^{\rm der}(E)$. We also note that ${\rm Ker}(\mathcal{J})$ is stable under the center $Z(E)$, and hence $\langle G(E).v\rangle$ is contained in ${\rm Ker}(\mathcal{J})$. Now, by the irreducibility of $\Pi$, we get a contradiction to the assumption that the map $\Phi$ is non-zero. ∎
@@ 3.6. Irreducibility of Tate cohomology for ${\rm GL}_{n}$
Let $G$ be the group ${\rm GL}_{n}$, and we prove the irreducibility of the ${\rm Gal}(E/F)$-Tate cohomology of $G(E)\rtimes{\rm Gal}(E/F)$-stable lattices in certain integral generic $\mathcal{K}$-representations of ${\rm GL}_{n}(E)$. We assume that $n=\ell m$ for some integer $m\geq 1$. Let $(\Pi,V)$ be the representation
@@
where $\tau$ is an absolutely irreducible cuspidal $\mathcal{K}$-representation of ${\rm GL}_{m}(E)$ such that $\tau$ is not isomorphic to twisted representation $\tau^{\sigma}$. Note that $\Pi\simeq\Pi^{\sigma}$, which induces an action of ${\rm Gal}(E/F)$ on the space $V$ that is compatible with the ${\rm Gal}(E/F)$-action on ${\rm GL}_{n}(E)$. We further assume that $\tau$ is integral, i.e., there exists a ${\rm GL}_{m}(E)$ stable $\Lambda$-lattice in the underlying space of $\tau$. The representation $\Pi$ admits a ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice in $V$. We now prove the following result.
Theorem 3.6.
@@
To see the cuspidality of the Tate cohomology groups $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{L})$, consider the Jacquet module $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{L})_{N(F)}$, where $N$ is the unipotent radical of a proper parabolic subgroup $P$. Let $M$ be the standard Levi subgroup of $P$. Using Proposition 3.2, we have $M(F)$ equivariant injection
$\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{L})_{N(F)}\hookrightarrow\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{L}_{N(E)}).$
We show that the Tate cohomology group $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{L}_{N(E)})$ is trivial for $i\in\{0,1\}$. Let $W$ be a finite length integral $\ell$-adic $M(E)\rtimes{\rm Gal}(E/F)$ representation such that the Jordan–Holder factors of $W$, as a representation of $M(E)$, are multiplicity free. Assume that all irreducible $M(E)$ sub-representation of $W$ are not stable under the action of ${\rm Gal}(E/F)$. The Jacquet module $V_{N(E)}$ is a special case of $W$ and it is integral as $\mathcal{L}_{N(E)}$ is a $M(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice in $V_{N(E)}$ by [Dat05, Proposition 1.4]. We use induction on the length of $W$ to show that $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{W})$ is trivial, for all $M(E)\rtimes{\rm Gal}(E/F)$-stable lattice $\mathcal{W}$ in $W$.
Let $W_{1}$ be an irreducible $M(E)$ sub-representation of $W$ and assume that $W_{1}$ is not fixed by ${\rm Gal}(E/F)$. The $\Lambda$-module
@@
is a $M(E)\rtimes{\rm Gal}(E/F)$ stable lattice in $W_{2}=\bigoplus_{i=0}^{l-1}\sigma^{i}(W_{1})$ and it is contained in the lattice $\mathcal{W}$. Since $\mathcal{M}$ is a projective $\Lambda[{\rm Gal}(E/F)]$-module, the Tate cohomology group $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{M})$ is trivial for $i\in\{0,1\}$. Now, using the long exact sequence of Tate cohomology corresponding to the following short exact sequence
@@
$\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{W})\simeq\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{W}/\mathcal{M}).$
Since the quotient $\mathcal{W}/\mathcal{M}$ is a $M(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice in $W/W_{2}$, and the length of $W/W_{2}$ is strictly less than the length of $W$, the Tate cohomology group $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{W}/\mathcal{M})$ is trivial (by induction). Therefore, $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{W})$ is trivial. Thus, we get that the space $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{L}_{N(E)})$ is trivial. Hence, the representation $\widehat{H}^{i}(\mathcal{L})$ is cuspidal. The $G(F)$ representation $\widehat{H}^{0}(\mathcal{L})$ admits a unique generic sub–quotient (Theorem 3.5). Since $\widehat{H}^{0}(\mathcal{L})$ is also cuspidal, it is irreducible. From the inclusion
$\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L})_{N(F),\Theta_{F}}\hookrightarrow\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L}_{N(E),\Theta_{E}}).$
we get that $\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L}_{N(E),\Theta_{E}})$ is trivial, and hence $\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L})_{N(F),\Theta_{F}}$ is trivial. Since $\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L})$ is also cuspidal, we get that $\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L})$ is trivial. ∎
Remark 3.7.
If $\Pi$ is an absolutely irreducible, integral, generic $\mathcal{K}$-representation of the form (3.6), and the $-\ell$ reduction $r_{\ell}(\Pi)$ is irreducible, then it follows from a similar argument as in Theorem 3.6 that $\widehat{H}^{0}(r_{\ell}(\Pi))$ is an irreducible cuspidal $\ell$-modular representation of ${\rm GL}_{n}(F)$.
@@ 4.1. Character formula for finite groups of Lie type
4.1.1.
We begin with the character theory of the finite group ${\rm GL}_{n}(\mathbb{F}_{q})$. We find it convenient to use Deligne–Lusztig theory to describe these characters here. Let us consider a connected reductive group $G$ defined over $\mathbb{F}_{q}$, and let $F:G\rightarrow G$ be the corresponding Frobenius endomorphism. Let $(S_{0},B_{0})$ be a pair consisting of an $F$-stable Borel subgroup $B_{0}$ and an $F$-stable maximal torus $S_{0}$ of $B_{0}$. The automorphism induced by the Frobenius endomorphism $F$ on the Weyl group $W(G,S_{0})$ is again denoted by $F$. For an $F$-stable maximal torus $T$ of $G$, there exists $w\in W(G,S_{0})$ such that $(T,F)$ is isomorphic to $(S_{0},{\rm Ad}(w)\circ F)$. Two $F$-stable maximal tori $T_{1}$ and $T_{2}$ of $G$ are conjugate in $G^{F}$ if and only if the corresponding Weyl group elements $w_{1}$ and $w_{2}$ are $F$-conjugate in $W(G,S_{0})$, i.e., if $w_{1}=ww_{2}F(w)^{-1}$, for some $w\in W(G,S_{0})$. For $w\in W(G,S_{0})$, we denote by $T_{w}$ the $F$-stable maximal torus of $G$ such that $(T_{w},F)$ is isomorphic to $(S_{0},{\rm Ad}(w)\circ F)$. For the above results, we refer to [DL76, Corollary 1.14].
4.1.2.
For an $F$-stable maximal torus $T$ and a character $\theta$ of $T^{F}$, Deligne–Lusztig associate a virtual representation $R^{G}_{T,\theta}$. Every irreducible $\ell$-adic representation of $G^{F}$ occurs in a virtual character $R^{G}_{T,\theta}$, for some torus $T$ and $\theta\in\widehat{T^{F}}$. For all unipotent elements $u\in G^{F}$, we denote by $Q_{G,T,F}(u)$, the Green’s function, which is defined as the value $R^{G}_{T,1}(u)$. Let $g\in G^{F}$ be an element with Jordan decomposition $g=su$. Let $C_{G}(s)^{0}$ be the connected component of the centralizer of $s\in G^{F}$. Let $T_{1},T_{2},\dots,T_{k}$ be representatives of the $(C_{G}(s)^{0})^{F}$-conjugacy classes of $F$-stable maximal tori in $C_{G}(s)^{0}$, which are conjugate to $T$ in $G^{F}$. For simplicity, we denote by $H$ the group $C_{G}(s)^{0}$. The Deligne–Lusztig character formula [DL76, Theorem 4.2] gives
@@
4.1.3.
When $G$ is ${\rm GL}_{n}$ and $F$ is the Frobenius map induced by $x\mapsto x^{q}$, the action of $F$ on $W(G,S_{0})\simeq S_{n}$ is trivial and the $F$-stable maximal tori of $G$ are parametrized by the conjugacy classes of $S_{n}$. Every cuspidal representation of $G^{F}$ is equal to $(-1)^{n-1}R_{T,\theta}$, where $T$ is the torus corresponding to the conjugacy class of $W(G,S_{0})$ containing $n$-cycles, and $\theta\in\widehat{T^{F}}$ is in general position, i.e., $\theta^{w}\neq\theta$ for all $w\in W(G,T)^{F}$. For an unipotent element $u\in G^{F}$, with an associated partition $\mu$ of the integer $n$, and a maximal torus corresponding to a partition $\lambda$ of the integer $n$, the Green’s function $Q_{G,T,F}(u)$ is explicity computed by J. A Green, and they are related to Hall polynomials. We will not make use of this relation to Hall polynomials, however, we will use the connection of Green’s polynomials with the representation theory of $S_{n}$ due to T. Springer. In the fundamental work ([Spr76, Theorem 7.8]), T. Springer showed that
@@
4.1.4.
Let $w\in S_{n}$ be an $n$-cycle and let $T_{w}$ be a $F$-stable maximal torus of ${\rm GL}_{n}$ corresponding to the element $w$. Let $s\in T^{F}$ be an element such that the characteristic polynomial of $s$ is $f(X)^{r}$ where $f(X)\in\mathbb{F}_{q}[X]$ is an irreducible polynomial of degree $d$. We have $n=dr$. Let $H$ be the connected group $C_{{\rm GL}_{n}}(s)$. Note that $H(\overline{\mathbb{F}}_{q})$ is isomorphic to ${\rm GL}_{r}(\overline{\mathbb{F}}_{q})^{d}$ and the Frobenius endomorphism $F:H\rightarrow H$ is given by
$(g_{1},g_{2},\dots,g_{d})\mapsto({\rm Fr}(g_{d}),{\rm Fr}(g_{1}),\dots,{\rm Fr}(g_{d-1})).$
Here ${\rm Fr}:{\rm GL}_{n}\rightarrow{\rm GL}_{n}$ is the Frobenius endomorphism of ${\rm GL}_{n}$ induced by the map $x\mapsto x^{q}$ on the entries of the matrices. Note that $H^{F}$ is isomorphic to ${\rm GL}_{r}(\mathbb{F}_{q^{d}})$. Let $D_{r}$ and $B_{r}$ be the diagonal torus and upper triangular Borel subgroup of ${\rm GL}_{r}$. Note that $D_{r}^{d}$ and $B_{r}^{d}$ give $F$-stable maximal torus and an $F$-stable Borel subgroup to be denoted by $S_{H}$ and $B_{H}$ respectively. The Frobenius endomorphism $F$ on $H$ induces a map $F^{\ast}$ on $W(H,S_{H})$. If we identify $W(H,S_{H})$ with $S_{r}^{d}$, then
@@
The next lemma is about $F^{\ell}$-stable maximal tori in $H$ which are conjugate to $T_{w}$ by an element in ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. Along the way we prove some standard observations which we could not find a convenient reference.
Lemma 4.1.
Let $w$ be an $n$-cycle in $S_{n}$ and let $T_{w}$ be the $F$-stable maximal torus of ${\rm GL}_{n}$. Let $s\in T_{w}^{F}$ be an element with characteristic polynomial $f(X)^{r}$ where $f(X)\in\mathbb{F}_{q}[X]$ is an irreducible polynomial of degree $d$.
@@
    Let $S$ be an $F$-stable maximal torus of $H$ such that $gSg^{-1}=T_{w}$ for some $g\in{\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. There exists a $h\in H^{F^{\ell}}$ such that $hSh^{-1}=T_{w}$.
@@
Let $w_{1}$ and $w_{2}$ be two elements of $W(H,S_{H})$ corresponding to $S_{1}$ and $S_{2}$ respectively. We identify the elements $w_{1}$ and $w_{2}$ by the $d$-touples $(s_{1},\dots,s_{d})$ and $(t_{1},\dots,t_{d})$ respectively, via the isomorphism
@@
where $s_{i},t_{j}\in S_{r}$ for $1\leq i,j\leq d$. Let $k\in S_{r}$ be an element such that

### v2

@@ 4. Some genericity and irreducibility results
An equivalent condition for $(\pi,V)$ to be $\mathfrak{m}_{F}$-generic–which we will use frequently–is that the twisted Jacquet module $V_{U(F),\Theta_{F}}\simeq R$. We also have the obvious notions of $\mathfrak{m}_{E}$-generic representations of $G(E)$. Let $(\pi,V)$ be a smooth generic $R$-representation of $G(F)$. Let $\mathcal{J}_{\pi}$ be a non-zero linear functional in ${\rm Hom}_{U(F)}(V,\Theta_{F})$. Let $\mathbb{W}(\pi,\Theta_{F})\subset{\rm Ind}_{U(F)}^{G(F)}(\Theta_{F})$ be the space of functions $W_{v}:G(F)\rightarrow R$, $v\in V$, where
@@
4.0.2.
We now prove that the ${\rm Gal}(E/F)$-Tate cohomology of a generic $\ell$-modular ${\rm Gal}(E/F)$ stable representation of $G(E)$ has a unique generic sub-quotient. We proved this result for the case where $G={\rm GL}_{n}/F$ using Kirillov model. However, we give a simpler proof here. Which will be used in the later sections to prove irreducibility of Tate cohomology groups of certain representations of ${\rm GL}_{n}$ obtained as Base change lifts of cuspidal representations.
Theorem 4.0.1.
Let $E/F$ be a finite Galois extension with $[E:F]=\ell$. Let $G$ be a split reductive algebraic group defined over $F$. Let $(\Pi,V)$ be a ${\rm Gal}(E/F)$ stable $\mathfrak{m}_{E}$-generic $\ell$-modular representation of $G(E)$. Then the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\Pi)$ admits a unique $\mathfrak{m}_{F}$-generic sub-quotient.
@@
Let $\Gamma$ be the Galois group ${\rm Gal}(E/F)$. Let $\mathcal{J}$ be a non-zero element in ${\rm Hom}_{U(E)}(\Pi,\Theta_{E})$. Let $\mathcal{W}$ be the Whittaker model of $\Pi$. Using Proposition 3.3.2, the Tate cohomology applied to the surjective map $\mathcal{W}\rightarrow\mathcal{W}_{U(E),\Theta_{E}}$ gives the following injection:
@@
Since ${\rm Gal}(E/F)$ acts trivially on the twisted Jacquet module $\mathcal{W}_{U(E),\Theta_{E}}\simeq\overline{\mathbb{F}}_{\ell}$, the Tate cohomology group $\widehat{H}^{0}(\mathcal{W}_{U(E),\Theta_{E}})$ is a one dimensional vector space over $\overline{\mathbb{F}}_{\ell}$. Now, let us consider the restriction to $G(F)$ map
@@
It is enough to show that $\Phi$ is non-zero. Suppose $\Phi(W_{v})=0$ for all $W_{v}\in\mathcal{W}^{\Gamma}$. Let $v\in{\rm Ker}(\mathcal{J})\cap\mathcal{W}^{\Gamma}$. Note that $\Pi(g)v\in\operatorname{Ker}(\mathcal{J})$, for $g\in G(F)$ and $g\in U(E)$. This shows that $\Pi(g)v\in{\rm Ker}(\mathcal{J})$, for all $g\in G^{\rm der}(E)$. Also note that ${\rm Ker}(\mathcal{J})$ is stable under the center $Z(E)$, and hence $\langle G(E).v\rangle$ is contained in ${\rm Ker}(\mathcal{J})$. Now, by the irreducibility of $\Pi$, we get a contradiction to the assumption that the map $\Phi$ is non-zero. ∎
@@ 4.1. Irreducibility of Tate cohomology for ${\rm GL}_{n}$
Let $G$ be the group ${\rm GL}_{n}$, and we prove the irreducibility of the ${\rm Gal}(E/F)$-Tate cohomology of ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$ stable lattices in certain integral generic $\mathcal{K}$-representations of ${\rm GL}_{n}(E)$. Assume that $n=\ell m$ for some integer $m\geq 1$. Let $(\Pi,V)$ be the irreducible representation
@@
where $\tau$ is an absolutely irreducible cuspidal $\mathcal{K}$-representation of ${\rm GL}_{m}(E)$ such that $\tau$ is not isomorphic to twisted representation $\tau^{\sigma}$. Note that $\Pi\simeq\Pi^{\sigma}$, which induces an action of ${\rm Gal}(E/F)$ on the space $V$ that is compatible with the ${\rm Gal}(E/F)$ action on ${\rm GL}_{n}(E)$. We further assume that $\tau$ is integral, i.e., there exists a ${\rm GL}_{m}(E)$ stable $\Lambda$-lattice in the underlying space of $\tau$. The representation $\Pi$ admits a ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice in $V$. We now prove the following result.
Theorem 4.1.1.
@@
We set $\Gamma={\rm Gal}(E/F)$. To see the cuspidality of the Tate cohomology groups $\widehat{H}^{i}(\Gamma,\mathcal{L})$, consider the Jacquet module $\widehat{H}^{i}(\Gamma,\mathcal{L})_{N(F)}$, where $N$ is the unipotent radical of an $F$-rational proper parabolic subgroup $P$ of ${\rm GL}_{n}$. Let $M$ be an $F$-rational Levi subgroup of $P$. Using Proposition 3.3.2, we have $M(F)$ equivariant injection
$\widehat{H}^{i}(\Gamma,\mathcal{L})_{N(F)}\hookrightarrow\widehat{H}^{i}(\Gamma,\mathcal{L}_{N(E)}).$
We show that the Tate cohomology group $\widehat{H}^{i}(\Gamma,\mathcal{L}_{N(E)})$ is trivial for $i\in\{0,1\}$. Say $W$ is a finite length integral $\ell$-adic $M(E)\rtimes\Gamma$ representation such that the Jordan–Holder factors of $W$, as a representation of $M(E)$, are multiplicity free. Assume that all irreducible $M(E)$ sub-representation of $W$ are not stable under the action of $\Gamma$. The Jacquet module $V_{N(E)}$ is a special case of $W$ and it is integral as $\mathcal{L}_{N(E)}$ is a $M(E)\rtimes\Gamma$ stable $\Lambda$-lattice in $V_{N(E)}$ by [Dat05, Proposition 1.4]. We use induction on the length of $W$ to show that $\widehat{H}^{i}(\Gamma,\mathcal{W})$ is trivial, for all $M(E)\rtimes\Gamma$ stable lattice $\mathcal{W}$ in $W$.
Let $W_{1}$ be an irreducible $M(E)$ sub-representation of $W$ and assume that $W_{1}$ is not fixed by $\Gamma$. The $\Lambda$-module
@@
is a $M(E)\rtimes\Gamma$ stable lattice in $W_{2}=\bigoplus_{i=0}^{\ell-1}\sigma^{i}(W_{1})$ and it is contained in the lattice $\mathcal{W}$. Since $\mathcal{M}$ is a projective $\Lambda[\Gamma]$-module, the Tate cohomology group $\widehat{H}^{i}(\Gamma,\mathcal{M})$ is trivial for $i\in\{0,1\}$. Now, using the long exact sequence of Tate cohomology corresponding to the following short exact sequence
@@
$\widehat{H}^{i}(\Gamma,\mathcal{W})\simeq\widehat{H}^{i}(\Gamma,\mathcal{W}/\mathcal{M}).$
Since the quotient $\mathcal{W}/\mathcal{M}$ is a $M(E)\rtimes\Gamma$ stable $\Lambda$-lattice in $W/W_{2}$, and the length of $W/W_{2}$ is strictly less than that of $W$, the Tate cohomology group $\widehat{H}^{i}(\Gamma,\mathcal{W}/\mathcal{M})$ is trivial (by induction). Therefore, $\widehat{H}^{i}(\Gamma,\mathcal{W})$ is trivial. Thus, we get that the Tate cohomology $\widehat{H}^{i}(\Gamma,\mathcal{L}_{N(E)})$ is trivial. Hence, the representation $\widehat{H}^{i}(\Gamma,\mathcal{L})$ is cuspidal. The $G(F)$ representation $\widehat{H}^{0}(\Gamma,\mathcal{L})$ admits a unique generic sub–quotient (Theorem 4.0.1). Since $\widehat{H}^{0}(\Gamma,\mathcal{L})$ is also cuspidal, it is irreducible. From the inclusion
$\widehat{H}^{1}(\Gamma,\mathcal{L})_{N(F),\Theta_{F}}\hookrightarrow\widehat{H}^{1}(\Gamma,\mathcal{L}_{N(E),\Theta_{E}}).$
we get that $\widehat{H}^{1}(\Gamma,\mathcal{L}_{N(E),\Theta_{E}})$ is trivial, and so is $\widehat{H}^{1}(\Gamma,\mathcal{L})_{N(F),\Theta_{F}}$. Since $\widehat{H}^{1}(\Gamma,\mathcal{L})$ is also cuspidal, we get that $\widehat{H}^{1}(\Gamma,\mathcal{L})$ is trivial. ∎
Remark 4.1.2.
If $\Pi$ is an absolutely irreducible, integral, generic $\mathcal{K}$-representation of the form (4.1), and the $-\ell$ reduction $r_{\ell}(\Pi)$ is irreducible, then it follows from a similar argument as in Theorem 4.1.1 that $\widehat{H}^{0}({\rm Gal}(E/F),r_{\ell}(\Pi))$ is an irreducible $\ell$-modular cuspidal representation of ${\rm GL}_{n}(F)$.
@@ 5.1. Character formula for finite groups of Lie type
5.1.1.
[…]
5.1.2.
For an $F$-stable maximal torus $T$ and a character $\theta$ of $T^{F}$, Deligne–Lusztig associate a virtual representation $R^{G}_{T,\theta}$. Every irreducible $\ell$-adic representation of $G^{F}$ occurs in a virtual character $R^{G}_{T,\theta}$, for some torus $T$ and $\theta\in\widehat{T^{F}}$. For all unipotent elements $u\in G^{F}$, we denote by $Q_{G,T,F}(u)$, the Green’s function, which is defined as the value $R^{G}_{T,1}(u)$. Let $g\in G^{F}$ be an element with Jordan decomposition $g=su$. Let $C_{G}(s)^{0}$ be the connected component of the centralizer of $s\in G^{F}$. Let $T_{1},T_{2},\dots,T_{k}$ be representatives of the $(C_{G}(s)^{0})^{F}$-conjugacy classes of $F$-stable maximal tori in $C_{G}(s)^{0}$, which are conjugate to $T$ in $G^{F}$. For simplicity, we denote by $H$ the group $C_{G}(s)^{0}$. The Deligne–Lusztig character formula (see [DL76, Theorem 4.2] or [GM20, Lemma 2.2.23, Chapter 2]) gives
@@
5.1.3.
When $G$ is ${\rm GL}_{n}$ and $F$ is the Frobenius endomorphism induced by $x\mapsto x^{q}$, the action of $F$ on $W(G,S_{0})\simeq S_{n}$ is trivial and the $F$-stable maximal tori of $G$ are parametrised by the conjugacy classes of $S_{n}$. Every cuspidal representation of $G^{F}$ is equal to $(-1)^{n-1}R_{T,\theta}^{G}$, where $T$ is the torus corresponding to the conjugacy class of $W(G,S_{0})$ containing $n$-cycles, and $\theta\in\widehat{T^{F}}$ is in general position, i.e., $\theta^{w}\neq\theta$ for all $w\in W(G,T)^{F}$. For an unipotent element $u\in G^{F}$ with associated partition $\mu$ of the integer $n$, and a maximal torus corresponding to a partition $\lambda$ of the integer $n$, the Green’s function $Q_{G,T,F}(u)$ is explicitly computed by J. A. Green, and they are related to Hall polynomials. We will not make use of this relation with Hall polynomials, however, we will use the connection of Green’s polynomials with the representation theory of $S_{n}$ due to T. Springer. In the fundamental work ([Spr76, Theorem 7.8]), T. Springer showed that
@@
5.1.4.
Let $F:{\rm GL}_{n}\rightarrow{\rm GL}_{n}$ be the Frobenius endomorphism of ${\rm GL}_{n}$ induced by the map $x\mapsto x^{q}$ on the entries of the matrices. Let $w\in S_{n}$ be an $n$-cycle, and let $T_{w}$ be an $F$-stable maximal torus of ${\rm GL}_{n}$ corresponding to the element $w$. Let $s\in T_{w}^{F}$ be an element such that the characteristic polynomial of $s$ is $f(X)^{r}$ where $f(X)\in\mathbb{F}_{q}[X]$ is an irreducible polynomial of degree $d$. We have $n=dr$. Let $H$ be the connected group $C_{{\rm GL}_{n}}(s)$. Note that $H(\overline{\mathbb{F}}_{q})$ is isomorphic to ${\rm GL}_{r}(\overline{\mathbb{F}}_{q})^{d}$ and the Frobenius endomorphism $F:H\rightarrow H$ is given by
$(g_{1},g_{2},\dots,g_{d})\longmapsto({\rm Fr}(g_{d}),{\rm Fr}(g_{1}),\dots,{\rm Fr}(g_{d-1})),$
where ${\rm Fr}$ is the Frobenius endomorphism of ${\rm GL}_{r}$ sending a matrix $(g_{ij})\in{\rm GL}_{n}(\overline{\mathbb{F}}_{q})$ to $(g_{ij}^{q})$. Then $H^{F}$ is isomorphic to ${\rm GL}_{r}(\mathbb{F}_{q^{d}})$. Let $D_{r}$ and $B_{r}$ be the diagonal torus and the upper triangular Borel subgroup of ${\rm GL}_{r}$. Note that $D_{r}^{d}$ and $B_{r}^{d}$ give an $F$-stable maximal torus and an $F$-stable Borel subgroup to be denoted by $S_{H}$ and $B_{H}$, respectively. The Frobenius endomorphism $F$ on $H$ induces a map $F^{\ast}$ on $W(H,S_{H})$. If we identify $W(H,S_{H})$ with $S_{r}^{d}$, then
@@
The next lemma is about the $F^{\ell}$-stable maximal tori in $H$ which are conjugate to $T_{w}$ by an element in ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. Along the way, we prove some standard observations which we could not find a convenient reference.
Lemma 5.1.1.
Let $w$ be an $n$-cycle in $S_{n}$, and let $T_{w}$ be an $F$-stable maximal torus of ${\rm GL}_{n}$ as above. Let $s\in T_{w}^{F}$ be an element with characteristic polynomial $f(X)^{r}$, where $f(X)\in\mathbb{F}_{q}[X]$ is an irreducible polynomial of degree $d$.
@@
    Let $S$ be an $F$-stable maximal torus of $H$ such that $gSg^{-1}=T_{w}$ for some $g\in{\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. There exists an element $h\in H^{F^{\ell}}$ such that $hSh^{-1}=T_{w}$.
@@
Let $w_{1}$ and $w_{2}$ be two elements of $W(H,S_{H})$ corresponding to $S_{1}$ and $S_{2}$, respectively. We identify the elements $w_{1}$ and $w_{2}$ by the $d$-tuples $(s_{1},\dots,s_{d})$ and $(t_{1},\dots,t_{d})$ respectively, via the isomorphism
@@
where $s_{i},t_{j}\in S_{r}$ for $1\leq i,j\leq d$. If there exists an element $k\in S_{r}$ such that